# Null-Space Filtering for Data-Free Continual Model Merging: Preserving Transparency, Promoting Fidelity

### v1

@@ A.2 Proof of Corollary 1
Let $B=P(\rho-\tau)^{\top}(\rho-\tau)P\succeq 0$ and $C=X^{\top}X\succeq 0$. By von Neumann’s trace inequality, $\mathrm{trace}(BC)\leq\|C\|_{2}\,\mathrm{trace}(B)$. Since $\|C\|_{2}=\|X^{\top}X\|_{2}=\sigma_{1}^{2}(X)$ and
@@ B.1 Details of Dataset and Task Settings
To comprehensively examine continual model merging, we curated a collection of 20 diverse image classification benchmarks, spanning natural objects, remote sensing, medical imagery, and text-rendered datasets. This selection largely follows prior practice in multi-domain evaluation (Tang et al., 2025), while ensuring balanced inclusion of datasets with varying granularity (from binary recognition to hundreds of categories). Specifically, the benchmarks include: SUN397 (Xiao et al., 2010), Stanford Cars (Krause et al., 2013), RESISC45 (Cheng et al., 2017), EuroSAT (Helber et al., 2019), SVHN (Netzer et al., 2011), GTSRB (Stallkamp et al., 2012), MNIST (LeCun et al., 1998), DTD (Cimpoi et al., 2014), Flowers102 (Nilsback & Zisserman, 2008), PCAM (Veeling et al., 2018), FER2013 (Goodfellow et al., 2013), Oxford-IIIT Pet (Parkhi et al., 2012), STL-10 (Coates et al., 2011), CIFAR-100 and CIFAR-10 (Krizhevsky & Hinton, 2009), Food-101 (Bossard et al., 2014), Fashion-MNIST (Xiao et al., 2017), EMNIST (Cohen et al., 2017), KMNIST (Clanuwat et al., 2018), and Rendered SST-2 (Socher et al., 2013).
@@ B.3 Details of Baselines
  • WUDI-Merging. A data-free merging method that reduces task interference by enforcing orthogonality between task vectors and their residual components (Cheng et al., ). In the continual setting, each linear layer $l$ maintains two task vectors: the cumulative merged vector from previous tasks $\tau^{\text{merged},(l)}_{t-1}=\theta^{\text{merged},(l)}_{t-1}-\theta_{0}^{(l)}$ and the current task vector $\tau_{t}^{(l)}=\theta_{t}^{(l)}-\theta_{0}^{(l)}$. The merged vector $\tau_{m}^{(l)}$ is obtained by optimizing $\mathcal{L}=\sum_{i}\frac{1}{\|\tau_{i,l}\|^{2}_{F}}\big\|(\tau_{m,l}-\tau_{i,l})(\tau_{i,l})^{\top}\big\|_{F}^{2}$. The update follows gradient descent, and the final merged parameters are $\theta^{\text{merged}}_{t}=\theta_{0}+\tau_{m}$.
@@ C.1 Detailed Overall Performance Results
Tab. 9 expands on the average results in Tab. 1 by reporting per-task average accuracy after continually merging 20 tasks. We compare six methods, SWA, Task Arithmetic, Ties-Merging, MagMax-IND, OPCM, and our proposed NUFILT across three CLIP-ViT backbones (B/32, B/16, L/14). NUFILT achieves the highest accuracy on most tasks. These fine-grained results reinforce the main paper’s findings, highlighting NUFILT’s ability to improve performance on continual model merging.
@@ C.2 Extended Visualizations on Subspace Alignment
This section provides extended subspace alignment visualizations for three additional models (ViT-B/32, ViT-L/14, Flan-T5-base; see Fig. 5, 6, and 7) to further validate the universal alignment of task vectors with task-relevant representations across architectures and layers. All visualizations follow a unified 3×3 grid structure where columns represent model components (left: full model; middle: attention layers; right: MLP layers) and rows show visualization types (top: layer-wise mean affinity heatmaps; middle: layer-wise 90th percentile affinity heatmaps; bottom: layer-wise ECDF curves). The layer-wise ECDFs reveal consistent spectral overlaps: for ViT models (Fig. 5 and 6), high (MNIST Data →MNIST Vector), moderate (EuroSAT Data →RESISC45 Vector), and low (SVHN Data →DTD Vector); for Flan-T5-base (Fig. 7), high (RTE Data →RTE Vector), moderate (RTE Data →QNLI Vector), and low (RTE Data →SST-2 Vector). These results across both vision and NLP domains confirm the consistent alignment phenomenon originally observed in Fig. 2, with all affinity heatmaps showing strong diagonal dominance where matched pairs exhibit significantly higher affinity than mismatched ones.
@@
Appendix D Discussions
D.1 Limitations
Similar to prior approaches in model merging, our NUFILT framework is built on the assumption that all task-specific fine-tuned models $\{\theta_{t}\}_{t=1}^{T}$ originate from a common pre-trained initialization $\theta_{0}$. The implications of this assumption—such as its effect on the alignment of task vectors in the underlying subspace—remain insufficiently explored and merit further investigation. Our current experiments are restricted to models sharing the same backbone (e.g., CLIP ViT variants, Flan-T5-base); extending the framework to heterogeneous initializations or architectures represents an interesting avenue for future work. Moreover, due to computational limitations, our study is confined to models with fewer than one billion parameters.
D.2 Broader Impacts
D.3 LLM Usage

### v2

@@ A.2 Proof of Corollary 1
Let $B=P(\rho-\tau)^{\top}(\rho-\tau)P\succeq 0$ and $C=X^{\top}X\succeq 0$. Since $C\preceq\|C\|_{2}I$, it follows that $\mathrm{trace}(BC)\leq\|C\|_{2}\,\mathrm{trace}(B)$. Since $\|C\|_{2}=\|X^{\top}X\|_{2}=\sigma_{1}^{2}(X)$ and
@@ B.1 Details of Dataset and Task Settings
To comprehensively examine continual model merging, we curated a collection of 20 diverse image classification benchmarks, spanning natural objects, remote sensing, medical imagery, and text-rendered datasets. This selection largely follows prior practice in multi-domain evaluation (Tang et al., 2025), while ensuring balanced inclusion of datasets with varying granularity (from binary recognition to hundreds of categories). Specifically, the benchmarks include: SUN397 (Xiao et al., 2010), Stanford Cars (Krause et al., 2013), RESISC45 (Cheng et al., 2017), EuroSAT (Helber et al., 2019), SVHN (Netzer et al., 2011), GTSRB (Stallkamp et al., 2012), MNIST (LeCun et al., 1998), DTD (Cimpoi et al., 2014), Flowers102 (Nilsback and Zisserman, 2008), PCAM (Veeling et al., 2018), FER2013 (Goodfellow et al., 2013), Oxford-IIIT Pet (Parkhi et al., 2012), STL-10 (Coates et al., 2011), CIFAR-100 and CIFAR-10 (Krizhevsky and Hinton, 2009), Food-101 (Bossard et al., 2014), Fashion-MNIST (Xiao et al., 2017), EMNIST (Cohen et al., 2017), KMNIST (Clanuwat et al., 2018), and Rendered SST-2 (Socher et al., 2013).
@@ B.3 Details of Baselines
  • WUDI-Merging. A data-free merging method that reduces task interference by enforcing orthogonality between task vectors and their residual components (Cheng et al., 2025). In the continual setting, each linear layer $l$ maintains two task vectors: the cumulative merged vector from previous tasks $\tau^{\text{merged},(l)}_{t-1}=\theta^{\text{merged},(l)}_{t-1}-\theta_{0}^{(l)}$ and the current task vector $\tau_{t}^{(l)}=\theta_{t}^{(l)}-\theta_{0}^{(l)}$. The merged vector $\tau_{m}^{(l)}$ is obtained by optimizing $\mathcal{L}=\sum_{i}\frac{1}{\|\tau_{i,l}\|^{2}_{F}}\big\|(\tau_{m,l}-\tau_{i,l})(\tau_{i,l})^{\top}\big\|_{F}^{2}$. The update follows gradient descent, and the final merged parameters are $\theta^{\text{merged}}_{t}=\theta_{0}+\tau_{m}$.
@@
B.4 Hyper-parameter Adjustment Guidelines
This section provides a brief guide for selecting the hyper-parameters used in NUFILT.
Rank hyper-parameters $(r_{p},r_{l},r_{v})$.
Our sensitivity analysis shows that all three rank hyper-parameters exhibit unimodal behavior and can be tuned independently with a small validation set:
  • $r_{p}$: Controls how much of the previous-task subspace is preserved. Larger values reduce forgetting; moderate values work consistently across backbones.
  • $r_{l}$: Low-rank correction term. Set smaller than $r_{p}$ to avoid disrupting the preserved subspace; small ranks are stable.
  • $r_{v}$: Dimensionality of the task vector subspace. A small value ensures stable alignment; increasing it introduces weaker singular directions and typically does not improve performance.
Number of Iterations.
Performance improves initially but saturates quickly. A single tuned value (e.g., 50–100 iterations) generalizes across different ViT backbones.
Scaling Across Architectures.
Sensitivity patterns are consistent for ViT-B/32, ViT-B/16, and ViT-L/14. The same hyperparameter settings work well across model sizes, with only minimal fine-tuning needed when desired.
@@ C.1 Detailed Overall Performance Results
Tab. 10 expands on the average results in Tab. 1 by reporting per-task average accuracy after continually merging 20 tasks. We compare six methods, WA, Task Arithmetic, Ties-Merging, MagMax-IND, OPCM, and our proposed NUFILT across three CLIP-ViT backbones (B/32, B/16, L/14). NUFILT achieves the highest accuracy on most tasks. These fine-grained results reinforce the main paper’s findings, highlighting NUFILT’s ability to improve performance on continual model merging.
@@ C.2 Extended Visualizations on Subspace Alignment
This section provides extended subspace alignment visualizations for three additional models (ViT-B/32, ViT-L/14, Flan-T5-base; see Fig. 6, 7, and 8) to further validate the universal alignment of task vectors with task-relevant representations across architectures and layers. All visualizations follow a unified 3×3 grid structure where columns represent model components (left: full model; middle: attention layers; right: MLP layers) and rows show visualization types (top: layer-wise mean affinity heatmaps; middle: layer-wise 90th percentile affinity heatmaps; bottom: layer-wise ECDF curves). The layer-wise ECDFs reveal consistent spectral overlaps: for ViT models (Fig. 6 and 7), high (MNIST Data →MNIST Vector), moderate (EuroSAT Data →RESISC45 Vector), and low (SVHN Data →DTD Vector); for Flan-T5-base (Fig. 8), high (RTE Data →RTE Vector), moderate (RTE Data →QNLI Vector), and low (RTE Data →SST-2 Vector). These results across both vision and NLP domains confirm the consistent alignment phenomenon originally observed in Fig. 2, with all affinity heatmaps showing strong diagonal dominance where matched pairs exhibit significantly higher affinity than mismatched ones.
Appendix D Discussions
D.1 Limitations
Similar to prior approaches in model merging, our NUFILT framework is built on the assumption that all task-specific fine-tuned models $\{\theta_{t}\}_{t=1}^{T}$ originate from a common pre-trained initialization $\theta_{0}$. The implications of this assumption—such as its effect on the alignment of task vectors in the underlying subspace—remain insufficiently explored and merit further investigation. Our current experiments are restricted to models sharing the same backbone (e.g., CLIP ViT variants, Flan-T5-base); extending the framework to heterogeneous initializations or architectures represents an interesting avenue for future work.
D.2 Broader Impacts
D.3 LLM Usage